OpenStudy (anonymous):

Strontium- 90 is used in machines that control the thickness of paper during the manufacturing process. Strontium-90 has a half life of 28 years. Determine how much time has elapsed if the following percentage of a sample remains at 3.125%.

OpenStudy (lgbasallote):

the formula for half-life is \[\frac{\ln 2}{k} = t\] where k is the decay constant and t is time \[\frac{\ln 2}{t} = k\] \[\frac{\ln 2}{28} = k\] now use the formula for decay \[\ln (\frac{x_o}{x}) = kt\] \[\Large \frac{\ln (\frac{x_o}{x})}{k} = t\] \[\Large \frac{\ln (\frac{x_o}{0.03125 x_o})}{\frac{\ln 2}{28}} = t\] hehe lol

OpenStudy (lgbasallote):

so basically \[\Large t = \frac{\ln (\frac{1}{0.03125})}{\frac{\ln 2}{28}}\]

OpenStudy (lgbasallote):

i'll let satellite do his traditional correction now...

OpenStudy (anonymous):

or maybe \[\left(\frac{1}{2}\right)^{\frac{t}{28}}=.0315\] \[\frac{t}{28}=\frac{\ln(.0315)}{\ln(.5)}\] \[t=\frac{28\ln(.0315)}{\ln(.5)}\]

OpenStudy (anonymous):

no just a different (and i find much easier) method since we are told the half life it is natural to use \(\frac{1}{2}\) as the base

OpenStudy (lgbasallote):

well yours definitely looks neater...

OpenStudy (lgbasallote):

lol yup they're the same :)